OpenStudy (anonymous):

i have a question, help me out.. The mean of four numbers is (p+2)^1/2. the sum of the squares of the numbers is 64 and the standard deviation is 2m. express p in terms of m

hartnn (hartnn):

know the formula for std deviation ?

OpenStudy (anonymous):

no

hartnn (hartnn):

\(\Large \sigma = \sqrt{\frac{1}{N} \sum \limits _{i=1}^N (x_i - \mu)^2}, {\rm \ \ where\ \ } \mu =Mean = \frac{1}{N} \sum \limits_{i=1}^N x_i.\)

hartnn (hartnn):

ever seen that ?

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

and then what?

hartnn (hartnn):

ok, so sigma = 2m N=4 x1 to x4 are our 4 numbers

OpenStudy (anonymous):

ok....

hartnn (hartnn):

from mean , we have x1+x2+x3+x4 = (p+2)^1/2 and x1^2+x2^2+x3^2+x4^2 =64 is given

OpenStudy (anonymous):

oh ok...

hartnn (hartnn):

nah, i made an error, (x1+x2+x3+x4)/4 = (p+2)^1/2

hartnn (hartnn):

let me check whether there's an easier alternate way to do this...

OpenStudy (anonymous):

hurm... i get p = 14 - 4m^2 is it correct

OpenStudy (anonymous):

what the easier alternate way to do this...

hartnn (hartnn):

how did you get that answer ?

hartnn (hartnn):

and yes i too get 4m^2 =14 -p

hartnn (hartnn):

which gives me p as 14 - 4m^2

hartnn (hartnn):

i did it the long way which i showed you

OpenStudy (anonymous):

how

hartnn (hartnn):

4m^2 =(1/4)[ (x1^2+x2^2+x^2+x4^2) -2 mu * (x1+x2+x3+x4) +4mu^2 )] if you plug in values and simplify, you'll easily get that

OpenStudy (anonymous):

oh, same as mine

hartnn (hartnn):

:)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^